### 21 Mar 2024 v5

# Instance-dependent uniform tail bounds for empirical processes

Abstract
We formulate a uniform tail bound for empirical processes indexed by a class of functions, in terms of the individual deviations of the functions rather than the worst-case deviation in the considered class. The tail bound is established by introducing an initial “deflation” step to the standard generic chaining argument. The resulting tail bound is the sum of the complexity of the “deflated function class” in terms of a generalization of Talagrand’s $\gamma$ functional, and the deviation of the function instance, both of which are formulated based on the natural seminorm induced by the corresponding Cramér functions. We also provide certain approximations for the mentioned seminorm when the function class lies in a given (exponential type) Orlicz space, that can be used to make the complexity term and the deviation term more explicit.
1 Introduction
Let $(X_{i})_{i=1}^{n}$ be i.i.d. copies of a random variable $X$ taking values in some space $\mathcal{X}$, and denote by $\operatorname{\mathbb{E}}_{n}$ the expectation with respect to the empirical measure associated with the samples $(X_{i})_{i=1}^{n}$. A central question of the theory of empirical processes is to find tail bounds for the empirical average $\operatorname{\mathbb{E}}_{n}f(X)=n^{-1}\sum_{i=1}^{n}f(X_{i})$ that hold uniformly for all functions $f$ belonging to a given function class $\mathcal{F}\subset{\mathbb{R}}^{\mathcal{X}}$.
Assuming that the functions in $\mathcal{F}$ are all zero-mean, the existing tail bounds in the literature typically assert that with probability at least $1-e^{-r}$, for all $f\in\mathcal{F}$ we have
$\operatorname{\mathbb{E}}_{n}f$ $\leq(\mathsf{C}(\mathcal{F})+\mathsf{S}_{r}(\mathcal{F}))o_{n}(1)\,,$
where $\mathsf{C}(\mathcal{F})$ depends on some measure of “complexity” of the function class $\mathcal{F}$ (e.g., VC-dimension [VC71, Vap98], Rademacher complexity [vdVW12, KP00], or Talagrand’s functional [Tal14]), and $\mathsf{S}_{r}(\mathcal{F})$ is some notion of the “worst-case deviation” of the functions $f\in\mathcal{F}$ at the confidence level $e^{-r}$. Our goal in this paper is to establish “instance-dependent” tail bounds in which the worst-case deviation above is replaced by the deviation of each particular function of the function class. It turns out that the instance-dependent tail bounds may provide some improvements in terms of the complexity term as well.
A closely related set of results are tail bounds for ratio type empirical processes. [GKW03, ] and [GK06, ] have developed such tail bounds for processes indexed by a class of $[0,1]$–bounded functions. In particular, various elaborate non-asymptotic tail bounds are derived in [GK06] by “slicing” the function class to sets of functions for which the variance proxy is nearly the same, and applying to each slice Talagrand’s concentration inequality for uniformly bounded empirical processes.
Our inspiration is a recent result of [LM23, ] on instance-dependent tail bounds for certain Gaussian processes. [LM23, ] used this result as a benchmark to motivate their main goal which is robust mean estimation with optimal direction-dependent sub-Gaussian confidence intervals. Specifically, in the case of a Gaussian processes indexed by points in some centered Euclidean ball, [LM23, Proposition 1] derived refined tail bounds that depend on the standard deviation at any queried direction rather than the worst-case standard deviation (i.e., largest eigenvalue of the corresponding covariance matrix). Furthermore, the complexity of the entire class is replaced by a quantity which, depending on the confidence level and the spectrum of the covariance matrix, can be significantly smaller than the square root of the trace of the covariance matrix appearing in the standard bounds.
Section 2 provides a more precise statement of the problem of interest. In Section Section 3 we consider two illustrative examples involving Gaussian processes. In particular, in Section 3.1, we discuss the problem studied by [LM23] in more details. Furthermore, in Section 3.2 we use instance-dependent bounds to formulate confidence intervals for the $m$th largest mean of a general Gaussian vector. The general instance-dependent tail bounds are presented in Section 4. Our calculations in Section 4.3 to derive more explicit expressions in the case of function classes in (exponential type) Orlicz spaces, can be of independent interest.
2 Preliminaries and Problem Setup
Let $\mathcal{F}$ denote a finite but arbitrarily large subset [^1] In many situations infinite function classes can be considered as well, but a completely rigorous analysis for the problems of interest requires the measurability issues to be addressed, e.g., as in [Pol84, Appx. C]. of a vector space $\mathbb{V}$ of centered functions from $\mathcal{X}$ to $\mathbb{R}$ whose cumulant generating function is finite in a neighborhood of the origin. Specifically, for every $f\in\mathbb{V}$ we have
$\operatorname{\mathbb{E}}f(X)$ $=0\,,$
and $I_{f}=\{\lambda\in\mathbb{R}\colon\log\operatorname{\mathbb{E}}e^{\lambda f(X)}<+\infty\}$, the domain of the corresponding cumulant generating function, contains 00 in its interior.
For simplicity we assume that the zero function, denoted by 00, is also in $\mathcal{F}$. We also frequently use functions $T_{r}\colon\,\mathbb{V}\to\mathbb{R}_{\geq 0}$ that are defined for $r\geq 0$ as
$T_{r}(g)$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\inf_{\lambda\geq 0}\frac{r+\log\operatorname{\mathbb{E}}e^{\lambda g(X)}}{\lambda}\,,$
with the convention that for $r>0$, if $\log\operatorname{\mathbb{E}}e^{\lambda g(X)}=+\infty$, then the objective of the infimum is also infinite and the corresponding $\lambda$ is implicitly excluded. These functions determine certain confidence intervals of interest and in fact are inverses of the rate function, a central object in the theory of large deviations [Var84, DZ10], associated with the random variable $g(X)$. We emphasize that the domain of $T_{r}(\cdot)$ is not restricted to $\mathcal{F}$, and as will be seen in the sequel we also apply $T_{r}(\cdot)$ to other functions in $\mathbb{V}$.
The function $T_{r}(\cdot)$ has certain properties that are important in our derivations. We have collected these properties in the following lemma that is proved in the Appendix.
Lemma 1 (Properties of $T_{r}(\cdot)$).
The function $T_{r}(\cdot)$ has the following properties:
  1. (i)
    $T_{r}(\cdot)$ is positive homogenous in the sense that $T_{r}(\alpha g)=\alpha T_{r}(g)$ for any $\alpha>0$ and all functions $g\in\mathbb{V}$.
  2. (ii)
    $T_{0}(g)=0$ for all functions $g$.
  3. (iii)
    The mapping $r\mapsto T_{r}(g)$, for $r>0$ and any particular function $g\in\mathbb{V}$, is concave and subadditive.
  4. (iv)
    The even envelope of $T_{r}(\cdot)$ defined as
    $\overline{T}_{r}(g)$ $=\max\{T_{r}(g),T_{r}(-g)\}\,,$ (1)
    is a seminorm.
For any fixed $f\in\mathbb{V}$ the following elementary lemma, which is essentially the well-known Chernoff bound, expresses a tail bound for $\mathbb{E}_{n}f$ in terms of $T_{r}(f)$. The proof is provided in the Appendix for completeness.
Lemma 2.
With the definitions above, for any function $f\in\mathbb{V}$ (whose moment generating function has 00 in the interior of its domain), with probability at least $1-e^{-nr}$, we have
$\operatorname{\mathbb{E}}_{n}f(X)$ $\leq T_{r}\left(f\right)\,.$
It is natural to seek an extension of Lemma 2 that provides an upper tail bound for the random variable of the form
$Z$ $=\sup_{f\in\mathcal{F}}\ \left(\operatorname{\mathbb{E}}_{n}f(X)-T_{r}\left(f\right)\right)\,,$
which translates to a uniform bound for $\operatorname{\mathbb{E}}_{n}f$ that holds for every instance of $f\in\mathcal{F}$. It is often more convenient to work with tail bounds expressed in terms of some seminorm of $f$ rather than the $T_{r}(f)$ which is not subadditive. A natural choice is $\overline{T}_{r}(\cdot)$ defined by (1), and consider the seminormed spaces $(\mathbb{V},\overline{T}_{r}(\cdot))$ for $r\geq 0$, where the functions of $\mathcal{F}$ belong to. In this paper we focus on finding an upper tail bound for random variables of the form
$Z$ $=\sup_{f\in\mathcal{F}}\ \left(\operatorname{\mathbb{E}}_{n}f(X)-\overline{T}_{r}(f)\right)\,.$
We emphasize that we use the term “instance-dependent tail bounds” specifically to refer to the bounds that generalize the Chernoff bound for an individual function, to the entire class as described above. For example, the result of the standard generic chaining arguments can be expressed in a way that the tail bounds depend on the queried function $f$. However, the resulting bounds are in terms of the optimal choice of the so-called admissible subsets of the function class, and the term $\overline{T}_{r}(f)$, even with a crude multiplicative factor, is not guaranteed to appear in the bound.
As an illustrative elementary example we derive a simple instance-dependent tail bound for empirical distribution of uniform random variables. In this example and in general, we may use certain accurate upper bounds for the terms $\overline{T}_{r}(f)$ which do not have an explicit expression, that are easier to evaluate.
Example 1.
Let $X$ be uniformly distributed over the interval $[0,1]$ and define $f_{t}(x)=\mathbb{1}_{\{x\leq t\}}-t$ for $x\in[0,1]$. We use the following bound for moment generating function of $f_{t}(X)$ as is commonly used in the derivation of the standard Bernstien’s inequality. For all $\lambda\in[-3,3]$, we have
$\operatorname{\mathbb{E}}e^{\lambda f_{t}(X)}$ $=1+\sum_{m=2}^{\infty}\frac{\operatorname{\mathbb{E}}f_{t}^{m}(X)}{m!}\lambda^{m}$
$\leq 1+\sum_{m=2}^{\infty}\frac{\operatorname{\mathbb{E}}f_{t}^{2}(X)}{m!}|\lambda|^{m}$
$\leq 1+\frac{t(1-t)}{2}\sum_{m=2}^{\infty}\lambda^{2}\left(\frac{|\lambda|}{3}\right)^{m-2}$
$\leq 1+\frac{t(1-t)}{2}\cdot\frac{\lambda^{2}}{1-|\lambda|/3}\,.$
Since $\log(1+u)\leq u$ for all $u>-1$, it follows that
$T_{r}(f_{t})$ $\leq\inf_{\lambda\in[0,3]}\frac{r+t(1-t)\lambda^{2}/(2-2\lambda/3)}{\lambda}$
$\leq\frac{1}{3}r+\sqrt{2t(1-t)r}\,,$
where the second line follows by evaluating the argument of the infimum at $\lambda=3r/(r+3\sqrt{t(1-t)/2})$. Therefore, it follows from (1) that $\overline{T}_{r}(f_{t})\leq r/3+\sqrt{2t(1-t)r}$. Our goal is then to find a bound for
$\sup_{t\in[0,1/2]}\operatorname{\mathbb{E}}_{n}f_{t}(X)-\frac{1}{3}(r+o_{n}(1))-\sqrt{2t(1-t)(r+o_{n}(1))}\,,$
that holds with high probability. [^2] Such a tail bound can be extended to the entire domain $t\in[0,1]$ using the symmetries of the problem. For $t\in[0,1/2]$ let $\widehat{t}=\lfloor(2n+1)t\rfloor/(2n+1)$ and decompose $f_{t}(x)$ as
$f_{t}(x)$ $=f_{t}(x)-f_{\widehat{t}}(x)+f_{\widehat{t}}(x)\,.$
Since $\widehat{t}$ takes only $n+1$ distinct values as $t$ varies from 00 to $1/2$, by a simple union bound, with probability at least $1-e^{-nr}$, for all $t\in[0,1/2]$ we have
$\operatorname{\mathbb{E}}_{n}f_{\widehat{t}}(X)-\frac{1}{3}(r+\log(n+1)/n)-\sqrt{2t(1-t)(r+\log(n+1)/n)}\leq 0\,,$ (2)
where we used the fact that $\widehat{t}(1-\widehat{t})\leq t(1-t)$ for every $t\in[0,1/2]$.
Furthermore, since $\sup_{t\in[0,1/2]}\operatorname{\mathbb{E}}(f_{t}(X)-f_{\widehat{t}}(X))^{2}<1/(2n+1)$, it follows from Bousquet’s inequality [Bou02, Theorem 2.3] that
$\sup_{t\in[0,1/2]}\operatorname{\mathbb{E}}_{n}\left(f_{t}(X)-f_{\widehat{t}}(X)\right)$ $\leq 2\operatorname{\mathbb{E}}\sup_{t\in[0,1/2]}\operatorname{\mathbb{E}}_{n}\left(f_{t}(X)-f_{\widehat{t}}(X)\right)+\sqrt{\frac{r}{n}}+\frac{4r}{3}\,,$ (3)
holds with probability at least $1-e^{-nr}$. With $(\varepsilon_{i})_{i\geq 1}$ denoting a sequence of Rademacher random variables independent of each other and everything else, by the symmetrization argument [BLM13, Lemma 11.4] we also have
$\operatorname{\mathbb{E}}\sup_{t\in[0,1/2]}\operatorname{\mathbb{E}}_{n}\left(f_{t}(X)-f_{\widehat{t}}(X)\right)$ $\leq 2\operatorname{\mathbb{E}}\sup_{t\in[0,1/2]}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\left(\mathbb{1}_{\{X_{i}\leq t\}}-\mathbb{1}_{\{X_{i}\leq\widehat{t}\}}\right)\right\rvert$
$\leq 2\operatorname{\mathbb{E}}\sup_{t\in[0,1/(2n+1)]}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\mathbb{1}_{\{X_{i}\leq t\}}\right\rvert\,.$ (4)
where the second line holds because the random variables $\mathbb{1}_{\{X\leq t\}}-\mathbb{1}_{\{X\leq\widehat{t}\}}$ and $\mathbb{1}_{\{X\leq t-\widehat{t}\}}$ have the same distribution, and $\widehat{t}\leq t<\widehat{t}+1/(2n+1)$. Let $N(t)$ denote the total number of indices $i$ such that $X_{i}\leq t$. Then, using the fact that $(\varepsilon_{i})_{i=1}^{n}$ are independent of $(X_{i})_{i=1}^{n}$, we deduce
$\operatorname{\mathbb{E}}\sup_{t\in[0,1/(2n+1)]}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\mathbb{1}_{\{X_{i}\leq t\}}\right\rvert$ $=\operatorname{\mathbb{E}}\left(\operatorname{\mathbb{E}}\left(\max_{j=0,1,2,\dotsc,N(1/(2n+1))}\left\lvert\frac{1}{n}\sum_{i=1}^{j}\varepsilon_{i}\right\rvert\Big{|}(X_{i})_{i=1}^{n}\right)\right)$
$\leq 2\operatorname{\mathbb{E}}\left(\left\lvert\frac{1}{n}\sum_{i=1}^{N(1/(2n+1))}\varepsilon_{i}\right\rvert\right)$
$\leq\frac{2}{n}\sqrt{\operatorname{\mathbb{E}}N(1/(2n+1))}$
$\leq\frac{\sqrt{2}}{n}\,,$ (5)
where the second line follows by writing the inner expectation as integration of the corresponding tail probabilities and invoking Lévy’s maximal inequality [BLM13, Lemma 11.12]. Combining (2), (3), (4), and (5) we conclude that with probability at least $1-2e^{-nr}$, for all $t\in[0,1/2]$ we have
$\operatorname{\mathbb{E}}_{n}f_{t}(X)$ $\leq\frac{1}{3}(r+\log(n+1)/n)+\sqrt{2t(1-t)(r+\log(n+1)/n)}$
$\quad+\frac{4\sqrt{2}}{n}+\sqrt{\frac{r}{n}}+\frac{4r}{3}\,.$
In particular, if we choose $r=\beta\log\log(n)/n$ for some $\beta>0$, and restrict $t$ to the interval $[1/n,1/2]$ so that $1/\sqrt{t}\leq\sqrt{n}$, we can normalize both sides of the derived instance-dependent tail bound by $\sqrt{t}$ and, with probability at least $1-2\log(n)^{-\beta}$, we have
$\sup_{t\in[1/n,1/2]}\frac{\operatorname{\mathbb{E}}_{n}f_{t}(X)}{\sqrt{t}}$ $\leq\frac{\beta\log\log(n)+\log(n+1)}{3\sqrt{n}}+\sqrt{\frac{2\beta\log\log(n)+2\log(n+1)}{n}}$
$\quad+\frac{4\sqrt{2}}{\sqrt{n}}+\sqrt{\frac{\beta\log\log(n)}{n}}+\frac{4\beta\log\log(n)}{3\sqrt{n}}\,.$ (6)
The derivations leading to (6) are not intended to produce the most refined bounds. In fact, it is known that the left-hand side of (6) is of order $\sqrt{\log\log(n)/n}$ (see, e.g., [Csá77], [GK06, Example 2.7], and [WS09, Chapter 16]).
3 Warm-up: Gaussian Processes
In this section we consider two examples involving Gaussian processes to further expose the structure and utility of instance-dependent bounds.
3.1 Marginals of a Gaussian Vector
Let us first consider a simple special case where the function class consists of linear functionals indexed by the centered unit Euclidean ball, i.e.,
=$\mathcal{F}=\{x\mapsto{\langle u,x\rangle}\colon\,u\in\mathbb{R}^{d}\,,{\left\lVert u\right\rVert}_{2}\leq 1\}\,,$
and the law of the underlying random variable $X\in\mathbb{R}^{d}$ is $\mathrm{Normal}(0,\Sigma)$. This scenario is studied in [LM23] who established the following instance-dependent tail bound.
Proposition 1 ([LM23, Proposition 1]).
For a nonnegative integer $k\leq d$, let $\Sigma_{k}$ denote the best rank-$k$ approximation of $\Sigma$ as measured by the operator norm. With probability at least $1-2e^{-nr}$, for all $u\in\mathbb{R}^{d}$ such that ${\left\lVert u\right\rVert}_{2}\leq 1$ we have
$\operatorname{\mathbb{E}}_{n}{\langle u,X\rangle}-\sqrt{2r}(u^{\scriptscriptstyle\textup{{T}}}\Sigma u)^{1/2}$ $\leq\sqrt{\frac{\operatorname{tr}(\Sigma-\Sigma_{k})}{n}}+\sqrt{{\left\lVert\Sigma-\Sigma_{k}\right\rVert}_{\mathrm{op}}}\,\sqrt{2r}+\sqrt{\frac{k}{n}}(u^{\scriptscriptstyle\textup{{T}}}\Sigma u)^{1/2}\,,$
where tr()normal-$\operatorname{tr}(\cdot)$ and ${\left\lVert\cdot\right\rVert}_{\mathrm{op}}$, respectively, denote the trace and the operator norm of their matrix arguments.
To understand the significance of Proposition 1 as well as the role of the integer parameter $k$, it is worth comparing the derived instance-dependent bound to the conventional bounds. A standard approach to bound $\operatorname{\mathbb{E}}_{n}{\langle u,X\rangle}$ uniformly for ${\left\lVert u\right\rVert}_{2}\leq 1$ is to apply the Gaussian concentration inequality (see, e.g., [BLM13, Theorem 5.6]) to ${\left\lVert\operatorname{\mathbb{E}}_{n}X\right\rVert}_{2}$, which, with probability at least $1-e^{-nr}$, guarantees that
$\sup_{u\colon{\left\lVert u\right\rVert}_{2}\leq 1}\operatorname{\mathbb{E}}_{n}{\langle u,X\rangle}={\left\lVert\operatorname{\mathbb{E}}_{n}X\right\rVert}_{2}\leq\sqrt{\frac{\operatorname{tr}(\Sigma)}{n}}+\sqrt{{\left\lVert\Sigma\right\rVert}_{\mathrm{op}}}\sqrt{2r}\,.$
This bound pessimistically considers the worst-case deviation for all of the random variables $\operatorname{\mathbb{E}}_{n}{\langle u,X\rangle}$. By setting $k=0$ in Proposition 1, we can reproduce this pessimistic bound, except for an extra factor of $2$ in front of the $r$-dependent term. A much better choice for $k$ in the instance-dependent tail bound can be found as follows. For $\ell=0,1,\dotsc,d$, let $d_{\ell}=\operatorname{tr}(\Sigma-\Sigma_{\ell})/{\left\lVert\Sigma-\Sigma_{\ell}\right\rVert}_{\mathrm{op}}$ denote the effective rank of $\Sigma-\Sigma_{\ell}$, and define
$k_{\star}$ $=\operatorname*{argmin}_{\ell=0,1,\dotsc,d}\max\left\{\frac{2nr}{d_{\ell}},\frac{\ell}{2nr}\right\}\,,$
and
$C_{\star}$ $=\max\left\{\sqrt{\frac{2nr}{d_{k_{\star}}}},\sqrt{\frac{k_{\star}}{2nr}}\right\}\,.$
Then, setting $k=k_{\star}$ and straightforward manipulations of the tail bound in Proposition 1 yields the inequality
$\operatorname{\mathbb{E}}_{n}{\langle u,X\rangle}$ $\leq(C_{\star}+1)\left(\sqrt{\frac{\operatorname{tr}(\Sigma-\Sigma_{k_{\star}})}{n}}+\sqrt{2r}(u^{\scriptscriptstyle\textup{{T}}}\Sigma u)^{1/2}\right)\,.$
Since $nr$ determines the confidence level of the tail bound, $k_{\star}$ and $C_{\star}$ only depend on this confidence level and the spectral characteristics of $\Sigma$. A favorable situation occurs when $C_{\star}$ is a small constant, which requires both $k_{\star}$ and $d_{k_{\star}}$ to be proportional to $nr$.
We provide a slightly different proof of Proposition 1 that is more streamlined and makes the constant factors reasonably small and explicit. Our proof only invokes the Gaussian concentration inequality, whereas the original proof in [LM23] uses the Gaussian Poincaré inequality as well.
Proof of Proposition 1.
Let $G\sim\mathrm{Normal(0,I)}$ be a standard normal random vector. Clearly, $\operatorname{\mathbb{E}}_{n}{\langle u,X\rangle}={\langle\frac{1}{\sqrt{n}}\Sigma^{1/2}u,G\rangle}$ in distribution, where $\Sigma^{1/2}$ is the symmetric square root of the covariance matrix $\Sigma$. Let $\pi_{k}G$ denote the orthogonal projection of $G$ onto the range of $\Sigma_{k}$. We have
${\langle\frac{1}{\sqrt{n}}\Sigma^{1/2}u,G\rangle}-(u^{\scriptscriptstyle\textup{{T}}}\Sigma u)^{1/2}\sqrt{2r}$
$\leq{\langle\frac{1}{\sqrt{n}}\Sigma^{1/2}u,G-\pi_{k}G\rangle}+{\langle\frac{1}{\sqrt{n}}\Sigma^{1/2}u,\pi_{k}G\rangle}-(u^{\scriptscriptstyle\textup{{T}}}\Sigma_{k}u)^{1/2}\sqrt{2r}$
$\leq\frac{1}{\sqrt{n}}{\left\lVert\Sigma^{1/2}(G-\pi_{k}G)\right\rVert}_{2}+\sqrt{\frac{u^{\scriptscriptstyle\textup{{T}}}\Sigma_{k}u}{n}}\left({\left\lVert\pi_{k}G\right\rVert}_{2}-\sqrt{2nr}\right)\,,$
where the second line follows from the fact that $u^{\scriptscriptstyle\textup{{T}}}\Sigma_{k}u\leq u^{\scriptscriptstyle\textup{{T}}}\Sigma u$, and the third line follows from the Cauchy-Schwarz inequality applied to each of the inner products. Using the Gaussian concentration inequality, with probability at least $1-e^{-nr}$ we have
${\left\lVert\Sigma^{1/2}(G-\pi_{k}G)\right\rVert}_{2}$ $\leq\sqrt{\operatorname{tr}(\Sigma-\Sigma_{k})}+\sqrt{{\left\lVert\Sigma-\Sigma_{k}\right\rVert}_{\mathrm{op}}}\,\sqrt{2nr}\,,$
and similarly, with probability at least $1-e^{-nr}$,
${\left\lVert\pi_{k}G\right\rVert}_{2}$ $\leq\sqrt{k}+\sqrt{2nr}\,.$
The desired result follows by combining the three derived inequalities. ∎
To put this special case in the general perspective, observe that, with $\mathbb{V}$ being the set of linear functionals over $\mathbb{R}^{d}$, the function class $\mathcal{F}$ consists of functions $f(x)=f_{u}(x)={\langle u,x\rangle}$ with ${\left\lVert u\right\rVert}_{2}\leq 1$, and we have
$T_{r}(f)=\overline{T}_{r}(f)=\sqrt{2r}\sqrt{u^{\scriptscriptstyle\textup{{T}}}\Sigma u}=\sqrt{2r}{\left\lVert f\right\rVert}_{L_{2}(X)}\,.$
3.2 Confidence Intervals for the “Middle-Ranked” Means of Correlated Gaussians
In this subsection we derive confidence intervals for the $m$-th largest mean of correlated Gaussian random variables, as another example where instance-dependent tail bounds can be applied. Similar to Proposition 1, we use the Gaussian concentration inequality, as well as a bound on the expected supremum of canonical Gaussian processes over (symmetric) polytopes [Tal14, Proposition 2.4.16 and Theorem 2.4.18] (see also the discussion in [vHan18, Section 3.3]). These tools allow us to express the upper and lower bounds of the confidence interval in more explicit terms.
Our goal is to find an upper and lower bounds for the $m$-th largest entry of a parameter vector $\theta\in\mathbb{R}^{d}$ for $m=o(d)$. We are only given $\widehat{\theta}=\theta+X$, where $X$ is a zero-mean Gaussian random variable with covariance $\Sigma=\operatorname{\mathbb{E}}XX^{\scriptscriptstyle\textup{{T}}}$. We assume that $\Sigma$ is known, and, without loss of generality, it is full-rank. For any vector $v$ we denote by $v^{\downarrow}$ the vector of the entries of $v$ sorted in decreasing order. Therefore, the $m$-th largest entry of a vector $v$ can be expressed as $v^{\downarrow}_{m}$. Furthermore, for any subset $S$ of $[d]\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\{1,\dotsc,d\}$ let $v_{S}\in\mathbb{R}^{|S|}$ denote the restriction of $v$ to the entries indexed by $S$. We also use the shorthand $\Sigma_{S}=\operatorname{\mathbb{E}}X_{S}X_{S}^{\scriptscriptstyle\textup{{T}}}$, which is the same as $\Sigma$ restricted to the rows and columns in $S$. By $\binom{[d]}{\ell}$, we denote the set of subsets of $[d]$ of size $\ell$, and we write $\triangle^{\ell}$ to denote the unit simplex in $\mathbb{R}^{\ell}$
Perhaps the simplest approach for our problem is to use the inequality
$\left|\theta^{\downarrow}_{m}-\widehat{\theta}^{\downarrow}_{m}\right|\leq{\left\lVert\theta-\widehat{\theta}\right\rVert}_{\infty}={\left\lVert X\right\rVert}_{\infty}\,,$
that suggests a confidence interval centered at the plug-in estimator $\widehat{\theta}^{\downarrow}_{m}$ whose width is no less than $2{\left\lVert X\right\rVert}_{\infty}$. The Gaussian concentration inequality then guarantees that ${\left\lVert X\right\rVert}_{\infty}\leq\operatorname{\mathbb{E}}{\left\lVert X\right\rVert}_{\infty}+\sqrt{2r}\max_{i\in[d]}\Sigma^{1/2}_{i,i}$ with probability at least $1-e^{-r}$. Furthermore, we can bound $\operatorname{\mathbb{E}}{\left\lVert X\right\rVert}_{\infty}$, viewed as the expected supremum of a canonical Gaussian process over a (symmetric) polytope, using [Tal14, Proposition 2.4.16 and Theorem 2.4.18]. Denoting the $i$th largest diagonal entry of $\Sigma$ by $\Sigma^{\downarrow}_{i,i}$, for some constant $C>0$ we have
$\left|\theta^{\downarrow}_{m}-\widehat{\theta}^{\downarrow}_{m}\right|$ $\leq C\max_{i\in[d]}\sqrt{\Sigma^{\downarrow}_{i,i}\log(i+1)}+\sqrt{2r}\sqrt{\Sigma^{\downarrow}_{1,1}}\,.$ (7)
Using the instance-dependent uniform tail bounds, we establish a more refined confidence interval for $\theta^{\downarrow}_{m}$ in the following proposition.
Proposition 2.
Let $m=o(d)$ be a positive integer, $r\in\mathbb{R}_{\geq 0}$, and $k\leq\min\{r,m\}$ be a nonnegative integer. Furthermore, for any nonempty set $S\subseteq[d]$ denote by $\Sigma_{S,k}$ the best rank-$k$ approximation of $\Sigma_{S}$ with respect to the operator norm, and define the vector $\sigma=\sigma(S,k)$ such that $\sigma_{i}=\sqrt{(\Sigma_{S}-\Sigma_{S,k})_{i,i}}$ for $i\in[|S|]$. Then, defining
$\sigma^{*}(S,k)$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\max_{i\in[|S|]}\sigma^{\downarrow}_{i}\sqrt{\log(i+1)}\,,$
$Q_{S,\beta}(\vartheta)$ $=\max_{u\in\triangle^{|S|}}{\langle u,\vartheta\rangle}-\beta{\left\lVert u\right\rVert}_{\Sigma_{S}}\,,$
which implicitly depends on $\Sigma_{S}$, and
$\beta_{r,m,k}$ $=\sqrt{r+m+m\log(d/m)+k}\,,$
with probability at least $1-4e^{-r}$ we have
$\theta^{\downarrow}_{m}$ $\geq\min_{S\in\binom{[d]}{d-m+1}}Q_{S,\beta_{r,m,k}}(\widehat{\theta}_{S})-C\sigma^{*}(S,k)-\sqrt{2}\beta_{r,m,k}{\left\lVert\sigma(S,k)\right\rVert}_{\infty}\,,$ (8)
and
$\theta^{\downarrow}_{m}$ $\leq\max_{S\in\binom{[d]}{m}}-Q_{S,\beta_{r,m,k}}(-\widehat{\theta}_{S})+C\sigma^{*}(S,k)+\sqrt{2}\beta_{r,m,k}{\left\lVert\sigma(S,k)\right\rVert}_{\infty}\,.$ (9)
Proof.
We express $\theta^{\downarrow}_{m}$ in an equivalent min-max variational form as
$\theta^{\downarrow}_{m}$ $=\min_{S\in\binom{[d]}{d-m+1}}\max_{u\in\triangle^{d-m+1}}{\langle u,\theta_{S}\rangle}$
$=\min_{S\in\binom{[d]}{d-m+1}}\max_{u\in\triangle^{d-m+1}}{\langle u,\widehat{\theta}_{S}\rangle}-{\langle u,X_{S}\rangle}\,.$ (10)
For the prescribed nonnegative integer $k\leq r$, let $\pi_{S}$ and $\pi_{S}^{\perp}$, respectively, denote the orthogonal projections onto the range and the nullspace of $\Sigma_{S,k}$. Then, with $G\sim\mathrm{Normal}(0,I)$ we can write
${\langle u,\pi_{S}X_{S}\rangle}$ =$\overset{\mathrm{d.}}{=}{\langle\Sigma_{S}^{1/2}\pi_{S}u,\pi_{S}G\rangle}$
$\leq{\left\lVert\Sigma_{S}^{1/2}\pi_{S}u\right\rVert}_{2}{\left\lVert\pi_{S}G\right\rVert}_{2}$
$={\left\lVert u\right\rVert}_{\Sigma_{S,k}}{\left\lVert\pi_{S}G\right\rVert}_{2}$
$\leq{\left\lVert u\right\rVert}_{\Sigma_{S}}{\left\lVert\pi_{S}G\right\rVert}_{2}\,.$
By the Gaussian concentration inequality, with probability at least $1-e^{-r}$, we have
${\left\lVert\pi_{S}G\right\rVert}_{2}$ $\leq\sqrt{2(r+k)}\,,$
thereby, on the same event, for all $u\in\mathbb{R}^{d-m+1}$ we have
${\langle u,\pi_{S}X_{S}\rangle}$ $\leq{\left\lVert u\right\rVert}_{\Sigma_{S}}\sqrt{2(r+k)}\,.$ (11)
Furthermore, recalling the definition of $\sigma=\sigma(S,k)$, with probability at least $1-e^{-r}$ we have
${\langle u,\pi^{\perp}_{S}X_{S}\rangle}$ $\leq{\left\lVert u\right\rVert}_{1}{\left\lVert\pi_{S}^{\perp}X_{S}\right\rVert}_{\infty}$
$\leq{\left\lVert u\right\rVert}_{1}(\operatorname{\mathbb{E}}({\left\lVert\pi_{S}^{\perp}X_{S}\right\rVert}_{\infty})+\sqrt{2r}{\left\lVert\sigma(S,k)\right\rVert}_{\infty})$
$\leq{\left\lVert u\right\rVert}_{1}(C\sigma^{*}(S,k)+\sqrt{2r}{\left\lVert\sigma(S,k)\right\rVert}_{\infty})\,,$
where the second line follows from the Gaussian concentration inequality, and the third line follows from [Tal14, Proposition 2.4.16 and the remarks after Theorem 2.4.18] for some absolute constant $C>0$. Adding the derived inequalities, with probability at least $1-2e^{-r}$, for all $u\in\mathbb{R}^{d-m+1}$, we can guarantee
${\langle u,X_{S}\rangle}$ $={\langle u,\pi^{\perp}X_{S}\rangle}+{\langle u,\pi X_{S}\rangle}$
$\leq{\left\lVert u\right\rVert}_{1}(C\sigma^{*}(S,k)+\sqrt{2r}{\left\lVert\sigma(S,k)\right\rVert}_{\infty})+{\left\lVert u\right\rVert}_{\Sigma_{S}}\sqrt{2(r+k)}\,.$ (12)
Applying this bound in (10), for any fixed $S\in\binom{[d]}{d-m+1}$, with probability at least $1-2e^{-r}$, we have
$\max_{u\in\triangle^{d-m+1}}{\langle u,\theta_{S}\rangle}$ $\geq\max_{u\in\triangle^{d-m+1}}{\langle u,\widehat{\theta}_{S}\rangle}-(C\sigma^{*}(S,k)+\sqrt{2r}{\left\lVert\sigma(S,k)\right\rVert}_{\infty})-{\left\lVert u\right\rVert}_{\Sigma_{S}}\sqrt{2(r+k)}\,.$
To obtain a lower bound for $\theta^{\downarrow}_{m}$, we can choose $S$ to be the indices of the $d-m+1$ smallest entries of $\theta$. But to be truly agnostic to the choice of $\theta$, we need to invoke the union bound and minimize the lower bound with respect to $S\in\binom{[d]}{d-m+1}$, at the cost of increasing $r$ by $m+m\log(d/m)>\log\tbinom{d}{m-1}$. The resulting inequality is then
$\theta^{\downarrow}_{m}$ $\geq\min_{S\in\binom{[d]}{d-m+1}}\max_{u\in\triangle^{d-m+1}}\Big{(}{\langle u,\widehat{\theta}_{S}\rangle}-(C\sigma^{*}(S,k)+\sqrt{2(r+m+m\log(d/m))}{\left\lVert\sigma(S,k)\right\rVert}_{\infty})$
$-{\left\lVert u\right\rVert}_{\Sigma_{S}}\sqrt{2(r+m+m\log(d/m)+k)}\Big{)}\,,$
which, by identifying the expressions of $\beta_{r,m,k}$ and $Q_{S,\beta_{r,m,k}}(\cdot)$, is equivalent to (8). To establish the upper bound (9), observe that $\theta^{\downarrow}_{m}=-(-\theta)^{\downarrow}_{d-m+1}$, which allows us to reuse the inequalities above to derive an upper bounds for $\theta^{\downarrow}_{m}$ through the lower bound for $(-\theta)^{\downarrow}_{d-m+1}$. ∎
It is worth mentioning that for $k=0$, the left-hand side of (11) vanishes, thereby we can improve the inequality (12) to
${\langle u,X_{S}\rangle}$ $\leq{\left\lVert u\right\rVert}_{1}(C\sigma^{*}(S,0)+\sqrt{2r}{\left\lVert\sigma(S,0)\right\rVert}_{\infty})\,.$
Consequently, for $k=0$ the corresponding bounds are in fact
$\theta^{\downarrow}_{m}$ $\geq\min_{S\in\binom{[d]}{d-m+1}}\left(\max_{i\in S}\widehat{\theta}_{i}-C\sigma^{*}(S,0)-\sqrt{2(r+m+m\log(d/m))}{\left\lVert\sigma(S,0)\right\rVert}_{\infty}\right)\,,$
and
$\theta^{\downarrow}_{m}$ $\leq\max_{S\in\binom{[d]}{m}}\left(\min_{i\in S}\widehat{\theta}_{i}+C\sigma^{*}(S,0)+\sqrt{2(r+m+m\log(d/m))}{\left\lVert\sigma(S,0)\right\rVert}_{\infty}\right)\,.$
If in addition to the assumption $m=o(d)$, we also have $m\lesssim r$ (i.e., $m\leq cr$ for some fixed constant $c>0$), the bounds above reproduce (7) up to an extra logarithmic factor for the term $\sqrt{\Sigma^{\downarrow}_{1,1}}$.
We also have the following minimax lower bounds for estimating $\theta^{\downarrow}_{m}$.
Proposition 3.
For $\kappa\geq 0$, let $\Theta=\{\theta\in\mathbb{R}^{d}\colon{\left\lVert\Sigma^{-1/2}\theta\right\rVert}_{2}\leq\kappa\}$ be a compact domain of parameters. With $\delta_{m}\geq 0$ defined as
$\delta_{m}$ $=\sup_{\theta\in\Theta}\theta^{\downarrow}_{m}+\sup_{\eta\in\Theta}\eta^{\downarrow}_{d-m+1}\,.$
For any estimator $g(\widehat{\theta})$ of $\theta^{\downarrow}_{m}$ we have
$\sup_{\theta\in\Theta}\operatorname{\mathbb{E}}(g(\widehat{\theta})-\theta^{\downarrow}_{m})^{2}$ $\geq\frac{\delta_{m}^{2}}{8e\max\{1,2\kappa^{2}\}}\,.$ (13)
Furthermore, we have
$\sup_{\theta\in\Theta}\operatorname{\mathbb{P}}\left(|g(\widehat{\theta})-\theta^{\downarrow}_{m}|>\frac{\delta_{m}}{3\max\{1,\sqrt{2}\kappa\}}\right)$ $\geq\frac{1}{2e}\,.$
Proof.
Le Cam’s two point method [PW24, Theorem 31.1] (see also [Yu97, Lemma 1]) guarantees that
$\sup_{\theta\in\Theta}\operatorname{\mathbb{E}}(g(\widehat{\theta})-\theta^{\downarrow}_{m})^{2}$ $\geq\sup_{\theta,\eta\in\Theta}\frac{1}{4}(\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m})^{2}\left(1-D_{\mathrm{TV}}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}_{\eta})\right)\,,$
where $D_{\mathrm{TV}}(\cdot,\cdot)$ denotes the total variation distance, and $\operatorname{\mathbb{P}}_{\theta}=\mathrm{Normal}(\theta,\Sigma)$ and $\operatorname{\mathbb{P}}_{\eta}=\mathrm{Normal}(\eta,\Sigma)$. The “simplified” Bretagnolle–Huber inequality [Tsy08, Equation 2.25] (see also [Can23] for a broader context) guarantees that
$D_{\mathrm{TV}}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}_{\eta})$ $\leq 1-\frac{1}{2}e^{-(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)/2}\,,$ (14)
using which we obtain
$\sup_{\theta\in\Theta}\operatorname{\mathbb{E}}(g(\widehat{\theta})-\theta^{\downarrow}_{m})^{2}$ $\geq\sup_{\theta,\eta\in\Theta}\frac{1}{8}(\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m})^{2}e^{-(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)/2}$
$=\sup_{\theta,\eta\in\Theta}\sup_{b\in[0,1]}\frac{1}{8}(\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m})^{2}be^{-b(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)/2}$
$\geq\frac{1}{8e\max\{1,2\kappa^{2}\}}\sup_{\theta,\eta\in\Theta}(\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m})^{2}\,,$
where the second line follows from the fact that $\Theta$ is star-shaped, and the third line follows from the inequality $\max_{b\in[0,1]}be^{-bz}\geq e^{-1}/\max\{1,z\}$ for $z\geq 0$. We can derive (13) from this lower bound using the fact that
$\sup_{\theta,\eta\in\Theta}(\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m})^{2}$ $=\left(\sup_{\theta\in\Theta}\theta^{\downarrow}_{m}-\inf_{\eta\in\Theta}\eta^{\downarrow}_{m}\right)^{2}$
$=\left(\sup_{\theta\in\Theta}\theta^{\downarrow}_{m}+\sup_{\eta\in\Theta}\eta^{\downarrow}_{d-m+1}\right)^{2}\,,$
where the latter equation follows from the symmetry of the set $\Theta$, and the fact that $-\eta^{\downarrow}_{m}=(-\eta)^{\downarrow}_{d-m+1}$.
Furthermore, it follows from the definition of the total variation distance and (14) that for any $c\geq 0$ we have
$|\operatorname{\mathbb{P}}(|g(\widehat{\theta})-\theta^{\downarrow}_{m}|>c)-\operatorname{\mathbb{P}}(|g(\widehat{\eta})-\theta^{\downarrow}_{m}|>c)|$ $\leq 1-\frac{1}{2}e^{-(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)/2}\,.$
In particular, for any $c\leq|\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m}|/2$, together with the inequality
$\operatorname{\mathbb{P}}(|g(\widehat{\eta})-\theta^{\downarrow}_{m}|>\frac{1}{2}|\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m}|)$ $\geq\operatorname{\mathbb{P}}(|g(\widehat{\eta})-\eta^{\downarrow}_{m}|<\frac{1}{2}|\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m}|)\,,$
which follows from the triangle inequality, we obtain
$\operatorname{\mathbb{P}}(|g(\widehat{\theta})-\theta^{\downarrow}_{m}|\geq c)+\operatorname{\mathbb{P}}(|g(\widehat{\eta})-\eta^{\downarrow}_{m}|\geq c)$ $\geq\frac{1}{2}e^{-(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)/2}\,.$
Therefore, if there exists a pair $\theta,\eta\in\Theta$ such that $(\theta-\eta)^{\scriptscriptstyle\textup{{T}}}\Sigma^{-1}(\theta-\eta)\leq 2$ and $|\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m}|/2\geq c$, then
$\sup_{\theta\in\Theta}\operatorname{\mathbb{P}}(|g(\widehat{\theta})-\theta^{\downarrow}_{m}|\geq c)$ $\geq\frac{1}{2e}\,.$
The desired result follows by setting $c=\sup_{\theta,\eta\in\Theta}|\theta^{\downarrow}_{m}-\eta^{\downarrow}_{m}|/(3\max\{1,\sqrt{2}\kappa\})$ which meets the required conditions. ∎
Because of the complicated and implicit form of the expressions in (8) and (9), it is difficult to compare—in full generality—the width of the confidence interval provided by Proposition 2 and the minimax lower bound of Proposition 3. We only focus on the special case where $\Sigma$ is diagonal. Furthermore, for the sake of simpler calculations we use the lower bound
$Q_{S,\beta}(\widehat{\theta}_{S})$ $\geq\max_{i\in S}\left(\widehat{\theta}_{i}-\beta\sqrt{\Sigma_{i,i}}\right)\,.$
The width of the confidence interval expressed by (8) and (9), which we denote by $\Delta_{m}$, can be bounded as
$\Delta_{m}=\max_{S\in\binom{[d]}{m}\,,S^{\prime}\in\binom{[d]}{d-m+1}}$ $\Big{(}-Q_{S,\beta_{r,m,k}}(-\widehat{\theta}_{S})-Q_{S^{\prime},\beta_{r,m,k}}(\widehat{\theta}_{S^{\prime}})$
$+C(\sigma^{*}(S,k)+\sigma^{*}(S^{\prime},k))$
$+\sqrt{2}\beta_{r,m,k}({\left\lVert\sigma(S,k)\right\rVert}_{\infty}+{\left\lVert\sigma(S^{\prime},k)\right\rVert}_{\infty})\Big{)}$
$\leq\max_{S\in\binom{[d]}{m}\,,S^{\prime}\in\binom{[d]}{d-m+1}}\Big{(}\min_{i\in S\,,j\in S^{\prime}}$ $\widehat{\theta}_{i}-\widehat{\theta}_{j}+\beta_{r,m,k}(\sqrt{\Sigma_{i,i}}+\sqrt{\Sigma_{j,j}})$
$+C(\sigma^{*}(S,k)+\sigma^{*}(S^{\prime},k))$
$+\sqrt{2}\beta_{r,m,k}({\left\lVert\sigma(S,k)\right\rVert}_{\infty}+{\left\lVert\sigma(S^{\prime},k)\right\rVert}_{\infty})\Big{)}$
$\leq\max_{S\in\binom{[d]}{m}\,,S^{\prime}\in\binom{[d]}{d-m+1}}\Big{(}\min_{i\in S\cap S^{\prime}}$ $2\beta_{r,m,k}\sqrt{\Sigma_{i,i}}+C(\sigma^{*}(S,k)+\sigma^{*}(S^{\prime},k))$
$+\sqrt{2}\beta_{r,m,k}({\left\lVert\sigma(S,k)\right\rVert}_{\infty}+{\left\lVert\sigma(S^{\prime},k)\right\rVert}_{\infty})\Big{)}\,,$
where the second inequality holds because $|S\cap S^{\prime}|=|S|+|S^{\prime}|-|S\cup S^{\prime}|\geq 1$, and we can choose $i=j\in S\cap S^{\prime}$. Furthermore, we have the inequalities
$\max\left\{{\left\lVert\sigma(S,k)\right\rVert}_{\infty},{\left\lVert\sigma(S^{\prime},k)\right\rVert}_{\infty}\right\}$ $\leq\sqrt{\Sigma^{\downarrow}_{1,1}}\,,$
and
$\max\left\{\sigma^{*}(S,k),\sigma^{*}(S^{\prime},k)\right\}$ $\leq\max_{i\in[d]}\sqrt{\Sigma^{\downarrow}_{i+k,i+k}\log(i+1)}\,,$
using which we deduce
$\Delta_{m}$ $\leq 2C\max_{i\in[d-k]}\sqrt{\Sigma^{\downarrow}_{i+k,i+k}\log(i+1)}+(2+2\sqrt{2})\beta_{r,m,k}\sqrt{\Sigma^{\downarrow}_{1,1}}\,.$
With $\Theta$ defined as in Proposition 3 we have
$\sup_{\theta\in\Theta}\theta^{\downarrow}_{m}$ $=\left(\sum_{i=1}^{m}\frac{1}{\Sigma^{\downarrow}_{i,i}}\right)^{-1/2}\kappa\,,$
and
$\sup_{\eta\in\Theta}\eta^{\downarrow}_{d-m+1}=\left(\sum_{i=1}^{d-m+1}\frac{1}{\Sigma^{\downarrow}_{i,i}}\right)^{-1/2}\kappa\,.$
Therefore, Proposition 3 implies that any confidence interval for $\theta^{\downarrow}_{m}$ with coverage probability no less than $1-1/(2e)$, should have a width equal to $C_{\kappa}\left(\sum_{i=1}^{m}\frac{1}{\Sigma^{\downarrow}_{i,i}}\right)^{-1/2}$ for some constant $C_{\kappa}\geq 0$ that may depend on $\kappa$. In particular, for any $\theta\in\Theta$ we have
$\operatorname{\mathbb{P}}\left(\Delta_{m}>C_{\kappa}\left(\sum_{i=1}^{m}\frac{1}{\Sigma^{\downarrow}_{i,i}}\right)^{-1/2}\right)$ $\geq\frac{1}{2e}\,.$
Choosing $r=1+\log(8)\approx 3$, we have also shown that
$\operatorname{\mathbb{P}}\left(\Delta_{m}>2C\max_{i\in[d-k]}\sqrt{\Sigma^{\downarrow}_{i+k,i+k}\log(i+1)}+(2+2\sqrt{2})\beta_{r,m,k}\sqrt{\Sigma^{\downarrow}_{1,1}}\right)$ $\leq\frac{1}{2e}\,.$
Then, if we define
$p_{m}(\Sigma)$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\sum_{i=1}^{m}\frac{\Sigma^{\downarrow}_{1,1}}{\Sigma^{\downarrow}_{i,i}}\,,$
and
$q_{m,k}(\Sigma)$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\max_{i\in[d-k]}\frac{\Sigma^{\downarrow}_{i+k,i+k}}{\Sigma^{\downarrow}_{1,1}}\log(i+1)\,,$
then $\Delta_{m}$ is optimal up to a factor $\mathrm{polylog}(d)$, if $m$, $p_{m}(\Sigma)$, and $q_{m,k}(\Sigma)$ all bounded from above as $\mathrm{polylog}(d)$. Specifically, if $m$, $p_{m}(\Sigma)$, and $q_{m,k}(\Sigma)$ are all absolute constants, then $\Delta_{m}$ is optimal up to a constant factor.
4 General Tail Bounds
We basically follow the generic chaining argument [Tal14] with an initial deflation of the function class that enables us to achieve the instance-dependence we aimed for. Furthermore, we use a “truncated chain” in our derivations similar to the approach of [Dir15, Theorem 3.2 ], with the distinction that we derive the tail bounds directly without resorting to the polynomial moments as in [Dir15].
4.1 Generalized $\gamma$ functional
Let us define $\rho_{r}(g,h)=\overline{T}_{r}(g-h)$ as a distance between a pair of functions $g,h\in\mathbb{V}$. With notation overloading, we also denote the distance of a function $g\in\mathbb{V}$ to a set of functions $\mathcal{H}\subseteq\mathbb{V}$ by
$\rho_{r}(g,\mathcal{H})$ $=\inf_{h\in\mathcal{H}}\overline{T}_{r}(g-h)\,.$ (15)
Similar to the truncated variant of Talagrand’s $\gamma$ functionals introduced in [Dir15], for $\mathcal{A}\subseteq\mathbb{V}$, and ¯$\underline{\ell}\in\mathbb{Z}_{\geq 0}$ we define
$\gamma(\mathcal{A};\,r,\underline{\ell})$ $=\inf_{\left(\mathcal{A}_{i}\right)_{i\geq 0}}\sup_{a\in\mathcal{A}}\ \sum_{\ell\geq\underline{\ell}}\rho_{r+(r+1/n)2^{\ell-\underline{\ell}}}(a,\mathcal{A}_{\ell})\,,$ (16)
where the infimum is taken over an increasing admissible sequence $(\mathcal{A}_{i})_{i\geq 0}$ of the subsets of $\mathcal{A}$ with $|\mathcal{A}_{i}|\leq 2^{2^{i}}$ for $i\geq 1$, and $\mathcal{A}_{0}=\{0\}$. For $n=1$, ¯$\underline{\ell}\approx\log_{2}(r)$, and $\overline{T}_{r}(g)=r^{1/\alpha}{\left\lVert g\right\rVert}_{\psi_{\alpha}}$ with ${\left\lVert\cdot\right\rVert}_{\psi_{\alpha}}$ being a $\psi_{\alpha}$ Orlicz norm, defined below in Section 4.3, the $\gamma$ functional defined in (16) effectively reduces to the Talagrand’s (truncated) $\gamma_{\alpha}$ functional. For a set $\mathcal{A}$, the Talagrand’s $\gamma_{\alpha}$ functional with respect to the suitable $\rho$ is defined as
$\gamma_{\alpha}(\mathcal{A},\rho;\,\underline{\ell})$ $=\inf_{\left(\mathcal{A}_{i}\right)_{i\geq 0}}\sup_{a\in\mathcal{A}}\sum_{\ell\geq\underline{\ell}}^{\infty}2^{\ell/\alpha}\rho(a,\mathcal{A}_{\ell})\,,$
where the infimum is again taken with respect to a sequence of admissible sets $(\mathcal{A}_{i})_{i\geq 0}$. The importance of these types of functionals was first revealed by Talagrand’s majorizing measures theorem [Tal87], whose appellation is due to the following essentially equivalent definition of $\gamma_{\alpha}(\mathcal{A},\rho)=\gamma_{\alpha}(\mathcal{A},\rho;0)$:
$\gamma_{\alpha}(\mathcal{A},\rho)$ $=\inf_{\mu}\sup_{a\in\mathcal{A}}\int_{0}^{\infty}\left(\log\frac{1}{\mu\left(\{b\in\mathcal{A}\colon\rho(b,a)\leq\varepsilon\}\right)}\right)^{1/\alpha}\mathrm{d}\varepsilon\,,$
with the infimum taken over probability measures $\mu$ on $\mathcal{A}$ [Tal01]. The majorizing measures theorem confirms a conjecture due to [Fer75, ] that the expectation of the supremum of the centered Gaussian process indexed by $\mathcal{A}$, is equivalent to $\gamma_{2}(\mathcal{A},\rho)$ up to constant factors, with $\rho$ being the canonical semimetric induced by the Gaussian process.
Evaluating or even finding a good approximation for a $\gamma$ functional of a general set $\mathcal{A}$ can be challenging [Tal01, vHan18], and the only solution could be “guessing” an appropriate majorizing measure or an admissible sequence of subsets [Tal01]. By pulling the supremum into the summation in the definition of $\gamma_{\alpha}$ functional, the infimum over the admissible sets would be achieved with each $\mathcal{A}_{i}$ being a covering set of $\mathcal{A}$ of cardinality $2^{2^{i}}$. This approximation describes the Dudley’s (entropy) integral inequality (see, e.g., [Ver18, Theorem 8.1.3], [Dir15, equation 2.3]), i.e.,
$\gamma_{\alpha}(\mathcal{A},\rho)\lesssim_{\alpha}\int_{0}^{\infty}\left(\log N(\mathcal{A},\rho,\varepsilon)\right)^{1/\alpha}\mathrm{d}\varepsilon\,,$
where $N(\mathcal{A},\rho,\varepsilon)$ is the covering number of $\mathcal{A}$ with respect to $\rho$-balls of radius $\varepsilon$, and $\lesssim_{\alpha}$ is the usual inequality sign up to a (positive) constant factor depending only on $\alpha$. If accurate estimates of the covering numbers of $\mathcal{A}$ are available, approximations of $\gamma_{\alpha}$ through Dudley’s inequality are easy to compute. However, Dudley’s inequality may not deliver sufficiently sharp approximations (see, e.g., [vHan18, Section 3.1]). The notable approach of [vHan18, ] improves on Dudley’s inequality by replacing the entropy numbers of the entire set $\mathcal{A}$ by those of certain scale-dependent “thin” subsets of $\mathcal{A}$, imitating the multiscale form of $\gamma_{\alpha}$. These thin subsets are “smoothed projections” of $\mathcal{A}$ expressed by minimizers of interpolation of the base metric and a given nonnegative functional at different scales [vHan18, Section 2.1]. The resulting approximation of $\gamma_{\alpha}$ is shown to be sharp in several nontrivial examples where Dudley’s inequality yields rather loose approximations [vHan18, Section 3].
It is worth mentioning that the $\gamma$ functional defined by (16) applies in more general settings than the standard $\gamma_{\alpha}$ functionals thanks to the less restricted form of the dependence of the pseudometric $\rho_{r}(\cdot,\cdot)$ on the “resolution scale” $r$. If $\mathcal{A}$, the function class of interest, is inhomogeneous in the sense that it contains functions with significantly different tail behavior, such as the function class of Example 1, then the standard $\gamma_{\alpha}$ functionals might overestimate the size (or complexity) of $\mathcal{A}$. As an illustrative example, suppose that for some absolute constant $\eta>0$ we have $\overline{T}_{r}(f)\approx{\left\lVert f\right\rVert}_{L_{\infty}}r+\eta{\left\lVert f\right\rVert}_{L_{2}}\sqrt{r}$ for all $f\in\mathbb{V}$, where approximation is in a multiplicative sense, and $L_{\infty}$ and $L_{2}$ norms are defined with respect to the law of $X$. This form of dependence on the resolution scale cannot be reproduced by the $\gamma_{\alpha}$ functional or other similarly defined quantities where the resolution scale and the distance to an admissible set are decoupled. Measuring the distance with respect to the scale-insensitive norm ${\left\lVert f\right\rVert}=c_{\infty}{\left\lVert f\right\rVert}_{L_{\infty}}+c_{2}{\left\lVert f\right\rVert}_{L_{2}}$ for arbitrary absolute constants $c_{2},c_{\infty}\geq 0$, leads to a suboptimal upper bound $\overline{T}_{r}(f)\leq(c^{\prime}_{\infty}r+c^{\prime}_{2}\sqrt{r})(c_{\infty}{\left\lVert f\right\rVert}_{L_{\infty}}+c_{2}{\left\lVert f\right\rVert}_{L_{2}})$ with $c^{\prime}_{2},c^{\prime}_{\infty}>0$ being constants that may depend only on $\eta$.
4.2 Main result: Generic Chaining with a “Deflation” Step
The following theorem is our main result.
Theorem 1.
Let $A\colon\mathcal{F}\to\mathcal{F}$ be a mapping such that
$\overline{T}_{r+k/n}(A[f])$ $\leq\overline{T}_{r+k/n}(f)\,,\quad\text{for all}\ f\in\mathcal{F}\,,$
and
$|A[\mathcal{F}]|$ $\leq e^{k}\,,$
for some nonnegative integer $k$, where $A[\mathcal{F}]=\left\{A[f]\colon f\in\mathcal{F}\right\}$ denotes the range of $A[\cdot]$. Furthermore, denote the “deflation” of $\mathcal{F}$ induced by $A[\cdot]$ by
$\mathcal{A}$ =$=\left\{f-A[f]\colon f\in\mathcal{F}\right\}\,.$
Setting ¯$\underline{\ell}=\lfloor\log_{2}(nr/3)\rfloor$ for $r\geq\log(2)/n$, with probability at least $1-2e^{-nr}$, for all $f\in\mathcal{F}$ we have
$\operatorname{\mathbb{E}}_{n}f(X)-\overline{T}_{r+k/n}(f)$ $\leq 2\gamma(\mathcal{A};\,r,\underline{\ell})+\operatorname{rad}_{2r+1/n}(\mathcal{A})\,,$ (17)
where $\gamma(\mathcal{A};\,r,\underline{\ell})$ is defined as in (16), and $\operatorname{rad}_{s}(\mathcal{A})=\max_{a\in\mathcal{A}}\overline{T}_{s}(a)$ denotes the radius of $\mathcal{A}$ measured by the seminorm $T_{s}(\cdot)$. The bound can further be optimized with respect to the mapping $A[\cdot]$, which both $k$ and $\mathcal{A}$ depend on.
Let us pause here to make a few remarks about Theorem 1. First, the effectiveness of the deflation step becomes clear by observing that the result of the standard generic chaining argument can be reproduced by the possibly suboptimal choice of $A[f]=0$ for all $f\in\mathcal{F}$ in (17). The admissible sequence in a standard generic chaining argument must cover $\mathcal{F}$, whereas in our formulation the admissible sequence must cover $\mathcal{A}$, the deflated version of $\mathcal{F}$. In particular, a desirable situation occurs when we can choose $A[\cdot]$ with $k\ll nr$ such that $\gamma(\mathcal{A};\,r,\underline{\ell})$ is smaller than $\gamma(\mathcal{F};\,r,0)$, and $\overline{T}_{r+k/n}(f)$ is close to $\overline{T}_{r}(f)$. Second, the assumption that $A[\mathcal{F}]$ is finite, is not essential; as can be seen below in the proof, it suffices to guarantee that $A[f]\leq\overline{T}_{r+o_{n}(1)}(f)$ holds, with probability at least $\geq 1-e^{-nr}$, for all $f\in\mathcal{F}$. For example, in Proposition 1, this condition is shown to hold through the Gaussian concentration inequality. Third, the right-hand side of (17) is basically an upper bound for $\sup_{a\in\mathcal{A}}\operatorname{\mathbb{E}}_{n}a(X)$ that holds with probability at least $1-e^{-nr}$. There are a few techniques to derive such upper bounds other than the generic chaining technique that we considered (see, e.g., [AB07] for a shortlist of the different techniques). For instance, in Example 1 the corresponding upper bound is obtained instead using the Bousquet’s inequality combined with Levy’s maximal inequality. The generic chaining has the advantage that it applies under rather general conditions, and in the case of Gaussian processes (as in the example of Section 3) yields sharp bounds.
Proof of Theorem 1.
As in (16), let $(\mathcal{A}_{\ell})_{\ell\geq 0}$ be an increasing admissible sequence of subsets of $\mathcal{A}$ such that $\mathcal{A}_{0}=\{0\}$. Let $c>0$ denote a constant that we will specify later in the proof, and set $r_{\ell}=r+(r+c/n)2^{\ell-\underline{\ell}}$. Given $A[\cdot]$ and the sequence $(\mathcal{A}_{\ell})_{\ell\geq 0}$, we can decompose every $f\in\mathcal{F}$ as
$f$ $=A[f]+f-A[f]$
$=A[f]+A_{\underline{\ell}}[f]+\sum_{\ell\geq\underline{\ell}}(A_{\ell+1}[f]-A_{\ell}[f])\,,$
where $A_{\ell}[f]$ denotes a function in $\mathcal{A}_{\ell}$ that is closest to $f-A[f]$ with respect to the seminorm $\overline{T}_{r_{\ell}}(\cdot)$, i.e.,
$A_{\ell}[f]$ $=\operatorname*{argmin}_{a\in\mathcal{A}_{\ell}}\overline{T}_{r_{\ell}}(f-A[f]-a)\,.$
It follows from the above decomposition that
$\operatorname{\mathbb{E}}_{n}f(X)-\overline{T}_{r+k/n}(f)$ $=\operatorname{\mathbb{E}}_{n}A[f](X)-\overline{T}_{r+k/n}(f)+$ (18)
$\qquad\operatorname{\mathbb{E}}_{n}A_{\underline{\ell}}[f](X)+\sum_{\ell\geq\underline{\ell}}\operatorname{\mathbb{E}}_{n}(A_{\ell+1}[f]-A_{\ell}[f])(X)\,.$
Lemma 2 and a simple union bound guarantee that, with probability at least $1-|A[\mathcal{F}]|e^{-nr-k}\geq 1-e^{-nr}$, we have
$\operatorname{\mathbb{E}}_{n}A[f](X)$ $\leq\overline{T}_{r+k/n}(A[f])\,,\quad\text{for all}\ f\in\mathcal{F}\,.$ (19)
Similarly, with probability at least $1-2^{2^{\underline{\ell}}}e^{-nr_{\underline{\ell}}}$, we have
$\operatorname{\mathbb{E}}_{n}A_{\underline{\ell}}[f](X)$ $\leq\overline{T}_{r_{\underline{\ell}}}(A_{\underline{\ell}}[f])\,,\quad\text{for all}\ f\in\mathcal{F}\,.$ (20)
Furthermore, for each index $\ell\geq\underline{\ell}$ there are at most $|\mathcal{A}_{\ell+1}|\,|\mathcal{A}_{\ell}|\leq 2^{2^{\ell+1}}2^{2^{\ell}}\leq 2^{2^{\ell+2}}$ different functions $A_{\ell+1}[f]-A_{\ell}[f]$ as $f$ varies in $\mathcal{F}$. Applying Lemma 2 and the union bound again it follows that, with probability at least $1-2^{2^{\ell+2}}e^{-nr_{\ell}}$, we also have
$\operatorname{\mathbb{E}}_{n}\left(A_{\ell+1}-A_{\ell}\right)[f](X)$ $\leq\overline{T}_{r_{\ell}}(\left(A_{\ell+1}-A_{\ell}\right)[f])\quad\text{for all}\ f\in\mathcal{F}\,.$ (21)
Putting (19), (20), and (21) back in the decomposition (18), with probability at least $1-e^{-nr}-2^{2^{\underline{\ell}}}e^{-nr_{\underline{\ell}}}-\sum_{\ell\geq\underline{\ell}}2^{2^{\ell+2}}e^{-nr_{\ell}}$, we have
$\operatorname{\mathbb{E}}_{n}f(X)-\overline{T}_{r+k/n}(f)$
$\leq\overline{T}_{r+k/n}(A[f])-\overline{T}_{r+k/n}(f)+\overline{T}_{r_{\underline{\ell}}}(A_{\underline{\ell}}[f])+\sum_{\ell\geq\underline{\ell}}\overline{T}_{r_{\ell}}\left(A_{\ell+1}[f]-A_{\ell}[f]\right)$
$\leq\operatorname{rad}_{r_{\underline{\ell}}}(\mathcal{A})+\sum_{\ell\geq\underline{\ell}}\overline{T}_{r_{\ell}}\left(A_{\ell+1}[f]-A_{\ell}[f]\right)$
$\leq\operatorname{rad}_{r_{\underline{\ell}}}(\mathcal{A})+\sum_{\ell\geq\underline{\ell}}\overline{T}_{r_{\ell}}\left(f-A[f]-A_{\ell}[f]\right)+\overline{T}_{r_{\ell}}\left(f-A[f]-A_{\ell+1}[f]\right)$
$\leq\operatorname{rad}_{r_{\underline{\ell}}}(\mathcal{A})+2\sum_{\ell\geq\underline{\ell}}\overline{T}_{r_{\ell}}\left(f-A[f]-A_{\ell}[f]\right)\,,\quad\text{for all}\ f\in\mathcal{F}\,,$
where the second inequality follows from the assumption $\overline{T}_{r+k/n}(A[f])\leq\overline{T}_{r+k/n}(f)$ and the fact that $\sup_{f\in\mathcal{F}}\overline{T}_{r_{\underline{\ell}}}(A_{\underline{\ell}}[f])\leq\sup_{a\in\mathcal{A}}\overline{T}_{r_{\underline{\ell}}}(a)$, and the third and fourth inequalities respectively follow from part (iv) of Lemma 1 and the fact that $\overline{T}_{r}(f)$ inherits the monotonicity with respect to $r$ from $T_{r}(f)$. Recalling the definition (15), on the same event we can write
$\operatorname{\mathbb{E}}_{n}f(X)-\overline{T}_{r+k/n}(f)$ $\leq\operatorname{rad}_{r_{\underline{\ell}}}(\mathcal{A})+2\sum_{\ell\geq\underline{\ell}}\rho_{r_{\ell}}\left(f-A[f],\mathcal{A}_{\ell}\right)\,,\quad\text{for all}\ f\in\mathcal{F}$
$\leq\operatorname{rad}_{r_{\underline{\ell}}}(\mathcal{A})+2\sup_{a\in\mathcal{A}}\sum_{\ell\geq\underline{\ell}}\rho_{r_{\ell}}\left(a,\mathcal{A}_{\ell}\right)\,.$
To complete the proof, it suffices to show that for $c=\log(2)<1$, and the prescribed ¯$\underline{\ell}=\left\lfloor\log_{2}(nr/3)\right\rfloor$, we have $2^{2^{\underline{\ell}}}e^{-nr_{\underline{\ell}}}+\sum_{\ell\geq\underline{\ell}}2^{2^{\ell+2}}e^{-nr_{\ell}}\leq e^{-nr}$. The specific choices of $c$ and $\underline{\ell}$ ensures that for $\ell\geq\underline{\ell}$ we have
$2^{2^{\ell+2}}e^{-n(r_{\ell}-r)}$ $=\left(2^{2^{\underline{\ell}+2}}e^{-nr-c}\right)^{2^{\ell-\underline{\ell}}}$
$\leq 2^{-(2^{\ell-\underline{\ell}})}\,.$
Furthermore, we have
$2^{2^{\underline{\ell}}}e^{-n(r_{\underline{\ell}}-r)}$ $\leq\frac{1}{16}\,.$
The desired inequality for the tail probability then follows as
$2^{2^{\underline{\ell}}}e^{-nr_{\underline{\ell}}}+\sum_{\ell\geq\underline{\ell}}2^{2^{\ell+2}}e^{-nr_{\ell}}$ $=\left(2^{2^{\underline{\ell}}}e^{-n(r_{\underline{\ell}}-r)}+\sum_{\ell\geq\underline{\ell}}2^{2^{\ell+2}}e^{-n(r_{\ell}-r)}\right)e^{-nr}$
$\leq\left(\frac{1}{16}+\frac{1}{2}+\sum_{\ell>\underline{\ell}}2^{-2^{\ell-\underline{\ell}}}\right)e^{-nr}$
$<\left(\frac{9}{16}+\frac{1/4}{1-1/4}\right)e^{-nr}$
$<e^{-nr}\,.$
4.3 Specializing to Orlicz Spaces
The purpose of this subsection is to approximate $T_{r}(f)$ for $f\in\mathbb{V}$, in situations where $\mathbb{V}$ is an Orlicz space of exponential type. Orlicz spaces are one of the important function spaces studied in functional analysis and probability theory. These function spaces can be described by their corresponding Orlicz norms. For a convex increasing function $\psi\colon{}[0,\infty)\to[0,\infty)$ with $\psi(0)=0$ the $\psi$-Orlicz norm of a random variable $Y$ is defined as
${\left\lVert Y\right\rVert}_{\psi}$ =$\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\inf\left\{u>0\colon\operatorname{\mathbb{E}}\psi\left(\frac{|Y|}{u}\right)\leq 1\right\}\,.$
Special cases are the usual $p$-norms for $p\geq 1$, the sub-Gaussian norm, and the sub-exponential norm, respectively, corresponding to $\psi(t)=t^{p}$, $\psi(t)=e^{t^{2}}-1$, and $\psi(t)=e^{t}-1$. Other interesting cases are the Bernstein–Orlicz norm corresponding to
$\psi(t)$ $=e^{{(\sqrt{1+2Lt}-1)}^{2}/L^{2}}-1\,,$
for some parameter $L>0$, introduced by [vdGL12, ], as well as the Bennett–Orlicz norm corresponding to
$\psi(t)$ $=e^{2\left((1+Lt)\log(1+Lt)-Lt\right)/L^{2}}-1\,,$
for some parameter $L>0$, introduced by [Wel17, ].
To express the general bounds presented in Theorem 1 when the underlying metric of interest imposed on $\mathcal{F}$ is induced by an Orlicz $\psi$-norm, it suffices to bound $T_{r}(f)$ in terms of ${\left\lVert f\right\rVert}_{\psi}\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}{\left\lVert f(X)\right\rVert}_{\psi}$. The following simple lemma can provide such bounds.
Lemma 3.
For every $f\in\mathbb{V}$ we have
$T_{r}(f)$ $\leq\inf_{\lambda\geq 0}\frac{r+\log\left(1+\int_{0}^{\infty}2\lambda\left(e^{\lambda t}-1\right)/\left(\psi(t)+1\right)\mathrm{d}t\right)}{\lambda}\,{\left\lVert f\right\rVert}_{\psi}\,.$ (22)
Proof.
Without loss of generality we may assume $f\neq 0$. The inequality follows by bounding the moment generating function of the zero-mean random variable $Y=f(X)/{\left\lVert f(X)\right\rVert}_{\psi}$, which has a unit $\psi$-Orlicz norm, as
$\operatorname{\mathbb{E}}e^{\lambda Y}$ $=\operatorname{\mathbb{E}}\left(e^{\lambda Y}-\lambda Y\right)$
$\leq\operatorname{\mathbb{E}}\left(e^{\lambda|Y|}-\lambda|Y|\right)$
$=1+\int_{0}^{\infty}\lambda\operatorname{\mathbb{P}}\left(|Y|>t\right)\left(e^{\lambda t}-1\right)\mathrm{d}t$
$\leq 1+\int_{0}^{\infty}\lambda\left(\operatorname{\mathbb{E}}\psi(|Y|)+1\right)\left(e^{\lambda t}-1\right)/\left(\psi(t)+1\right)\mathrm{d}t$
$=1+\int_{0}^{\infty}2\lambda\left(e^{\lambda t}-1\right)/\left(\psi(t)+1\right)\mathrm{d}t\,.\qed$
For exponential type Orlicz norms, defined below, we have the following proposition that provides a more explicit approximation for $T_{r}(f)$ in terms of ${\left\lVert f\right\rVert}_{\psi}$.
Proposition 4.
Let ${\left\lVert\cdot\right\rVert}_{\psi}$ be an Orlicz norm of exponential type, meaning that
$\psi(t)$ $=e^{\phi(t)}-1\,$
for a convex and increasing function $\phi\colon{}[0,\infty)\to[0,\infty)$ with $\phi(0)=0$. Furthermore, let ϕ*()superscriptitalic-$\phi^{*}(\cdot)$ denote the convex conjugate of ϕ()italic-$\phi(\cdot)$, i.e.,
$\phi^{*}(\lambda)$ $=\sup_{t\geq 0}\left(\lambda t-\phi(t)\right)\,.$
If for some $M>0$ we have
$\inf_{\lambda\geq 0}\ \frac{e^{\phi^{*}(\lambda)}-1}{\lambda^{2}}$ $\geq M\int_{0}^{\infty}te^{-\phi(t)/2}\mathrm{d}t\,,$ (23)
then for every $f\in\mathbb{V}$ we have
$T_{r}(f)$ $\leq\max\{3,3/\sqrt{2M}\}\phi^{-1}(2r/3)\,{\left\lVert f\right\rVert}_{\psi}\,.$ (24)
Proof.
Since $\phi^{*}(\cdot)$ is the convex conjugate of $\phi(\cdot)$, for every $\lambda,t\geq 0$ we can write
$\lambda t\leq\frac{1}{2}\phi(t)+\frac{1}{2}\phi^{*}(2\lambda)\,.$
Applying this bound to (22) of Lemma 3 we have
$T_{r}(f)$ $\leq\inf_{\lambda\geq 0}\frac{r+\log\left(1+\int_{0}^{\infty}2\lambda\left(1-e^{-\lambda t}\right)e^{\lambda t-\phi(t)}\mathrm{d}t\right)}{\lambda}\,{\left\lVert f\right\rVert}_{\psi}$
$\leq\inf_{\lambda\geq 0}\frac{r+\log\left(1+2\lambda^{2}e^{\phi^{*}(2\lambda)/2}\int_{0}^{\infty}te^{-\phi(t)/2}\mathrm{d}t\right)}{\lambda}\,{\left\lVert f\right\rVert}_{\psi}\,,$
where we also used the inequality $1-e^{-\lambda t}\leq\lambda t$. It follows from (23) that
$2\lambda^{2}\int_{0}^{\infty}te^{-\phi(t)/2}\mathrm{d}t$ $\leq e^{\phi^{*}(\sqrt{2/M}\lambda)}-1\,.$
Then, using the fact that $\phi^{*}(\cdot)$ is nonnegative, we have
$T_{r}(f)$ $\leq\inf_{\lambda\geq 0}\frac{r+\phi^{*}(2\lambda)/2+\phi^{*}(\sqrt{2/M}\lambda)}{\lambda}\,{\left\lVert f\right\rVert}_{\psi}\,.$
Furthermore, because $\phi^{*}(\cdot)$ is increasing, we can write
$\phi^{*}(2\lambda)/2+\phi^{*}(\sqrt{2/M}\lambda)$ $\leq\frac{3}{2}\phi^{*}\left(\max\{2,\sqrt{2/M}\}\lambda\right)\,.$
Therefore, we conclude that
$T_{r}(f)$ $\leq\inf_{\lambda\geq 0}\frac{r+\frac{3}{2}\phi^{*}\left(\max\{2,\sqrt{2/M}\}\lambda\right)}{\lambda}\,{\left\lVert f\right\rVert}_{\psi}$
$=\max\{3,3/\sqrt{2M}\}\phi^{-1}(2r/3)\,{\left\lVert f\right\rVert}_{\psi}\,.\qed$
It is worth mentioning that the constants appearing in the proposition are not necessarily optimal. In fact, the result may be improved for example by using the bound $1-e^{-\lambda t}\leq\min\{\lambda t,1\}$ instead of the inequality $1-e^{-\lambda t}\leq\lambda t$ that is used in the current proof. We did not pursue these refinements intending to obtain relatively simpler expressions.
Let us quantify the result of Proposition 4 when ${\left\lVert\cdot\right\rVert}_{\psi}$ is the sub-Gaussian Orlicz norm, and when it is the Bernstein–Orlicz norm. In the sub-Gaussian case, we have $\phi(t)=t^{2}$ and $\phi^{*}(\lambda)=\mathbb{1}(\lambda\geq 0)\lambda^{2}/4$. It is easy to verify that (23) holds with $M=1/4$. Therefore, for the sub-Gaussian Orlicz norm, (24) reduces to
$T_{r}(f)$ $\leq\sqrt{12r}\,{\left\lVert f\right\rVert}_{\psi}\,.$
In the case of Bernstein–Orlicz norm, $\phi(t)=(\sqrt{1+2Lt}-1)^{2}/L^{2}$. By the change of variable $t=\left((Lu+1)^{2}-1\right)/(2L)$ and using standard Gaussian integral formulas we can calculate the integral on the right-hand side of (23) as
$\int_{0}^{\infty}te^{-\left(\sqrt{1+2Lt}-1\right)^{2}/(2L^{2})}\mathrm{d}t$ $=\sqrt{\frac{\pi}{8}}L+1\,.$
Furthermore, with some straightforward calculations we can show that the convex conjugate of $\phi(\cdot)$ is
$\phi^{*}(\lambda)$ =$=\begin{cases}0\,,&\lambda<0\,,\\ \frac{\lambda^{2}}{4\left(1-L\lambda/2\right)}\,,&\lambda\in[0,2/L)\,,\\ \infty\,,&\lambda>2/L\,.\end{cases}$
Therefore, for $\lambda\geq 0$, we have
$e^{\phi^{*}(\lambda)}-1$ $\geq\phi^{*}(\lambda)\geq\frac{\lambda^{2}}{4}\,.$
Consequently, (23) holds if
$M$ $=\frac{1}{\sqrt{2\pi}L+4}\,,$
for which (24) reduces to
$T_{r}(f)$ $\leq 3(\sqrt{\pi/2}L+2)^{1/2}\phi^{-1}(2r/3)\,{\left\lVert f\right\rVert}_{\psi}$
$=(\sqrt{\pi/2}L+2)^{1/2}\left(Lr+\sqrt{6r}\right)\,{\left\lVert f\right\rVert}_{\psi}\,.$
Remaining Proofs
Proof of Lemma 1.
Part (i) of the lemma follows from a straightforward change of variable.
For part (ii), we have
$T_{0}(f)$ $=\inf_{\lambda\geq 0}\frac{\log\operatorname{\mathbb{E}}e^{\lambda f(X)}}{\lambda}$
$\leq\lim_{\lambda\downarrow 0}\frac{\log\operatorname{\mathbb{E}}e^{\lambda f(X)}}{\lambda}$
$=\lim_{\lambda\downarrow 0}\frac{\operatorname{\mathbb{E}}\left(f(X)e^{\lambda f(X)}\right)}{\operatorname{\mathbb{E}}e^{\lambda f(X)}}$
$=0\,,$
where the third and fourth line respectively follow from the l’Hôpital’s rule and the assumption that $f(X)$ is zero-mean. However, by Jensen’s inequality we have $\operatorname{\mathbb{E}}e^{\lambda f(X)}\geq e^{\lambda\operatorname{\mathbb{E}}f(X)}=1$, which means that $T_{0}(f)\geq 0$. Therefore, we must have
$T_{0}(f)$ $=0\,.$
For part (iii) observe that $T_{r}(f)$ can be equivalently expressed as
$T_{r}(f)$ $=\inf_{\theta\geq 0}\left(\theta r+\theta\log\operatorname{\mathbb{E}}e^{f(X)/\theta}\right)$
$=-\sup_{\theta\geq 0}\left(-\theta r-\theta\log\operatorname{\mathbb{E}}e^{f(X)/\theta}\right)\,.$
Since the supremum is the convex conjugate of $\theta\mapsto\theta\log\operatorname{\mathbb{E}}e^{f(X)/\theta}$ evaluated at $-r$, we conclude that $r\mapsto T_{r}(f)$ is concave. The proved concavity together with part (ii) of the lemma, guarantee that for all $r,s\geq 0$ we have
$\frac{r}{r+s}T_{r+s}(f)$ $=\frac{r}{r+s}T_{r+s}(f)+\frac{s}{r+s}T_{0}(f)$
$\leq T_{r}(f)\,,$
and
$\frac{s}{r+s}T_{r+s}(f)$ $=\frac{s}{r+s}T_{r+s}(f)+\frac{r}{r+s}T_{0}(f)$
$\leq T_{s}(f)\,,$
which add up to
$T_{r+s}(f)$ $\leq T_{r}(f)+T_{s}(f)\,,$
proving the subadditivity of $r\mapsto T_{r}(f)$.
To prove part (iv) we readily have $\overline{T}_{r}(0)=0$, and
$\overline{T}_{r}(\alpha f)$ $=\max\{T_{r}(\alpha f),T_{r}(-\alpha f)\}$
$=|\alpha|\overline{T}_{r}(f)\,,$
for every $f\in\mathbb{V}$ and nonzero real number $\alpha$. Therefore, it suffices to show that $\overline{T}_{r}(f)$ is convex in $f\in\mathbb{V}$. We show that $T_{r}(\cdot)$ is convex, which implies the convexity of $\overline{T}_{r}(\cdot)$. Let us define $\kappa(f)=\log\operatorname{\mathbb{E}}e^{f(X)}$ for $f\in\mathbb{V}$, and denote by $\mathbb{V}^{*}$ the dual space of $\mathbb{V}$, i.e., the space of linear functionals on $\mathbb{V}$ that are bounded in the sup norm. It follows from the Hölder’s inequality that $\kappa(\cdot)$ is convex. We also define the convex conjugate of $\kappa(\cdot)$ as
$\kappa^{*}(w)=\sup_{f\in\mathbb{V}}{\langle w,f\rangle}-\kappa(f)\,,$
for every $w\in\mathbb{V}^{*}$. It can be shown that $\kappa(\cdot)$ is also lower semi-continuous which guarantees $\kappa(f)=\sup_{w\in\mathbb{V}^{*}}{\langle w,f\rangle}-\kappa^{*}(w)$ for all $f\in\mathbb{V}$. Our goal is to show that
$T_{r}(f)$ $=\sup_{w\in\mathbb{V}^{*}\colon\kappa^{*}(w)\leq r}{\langle w,f\rangle}\,,$ (25)
which clearly proves the convexity of $T_{r}(f)$. For $r=0$ the identity (25) holds trivially as $\kappa(f)=0$ for all $f\in\mathbb{V}$. Then, without loss of generality we may assume that $r>0$ and write the right-hand side of (25) as
$\sup_{w\in\mathbb{V}^{*}\colon\kappa^{*}(w)\leq r}{\langle w,f\rangle}$ $=\sup_{w\in\mathbb{V}^{*}}\inf_{\gamma\geq 0}{\langle w,f\rangle}-\gamma(\kappa^{*}(w)-r)\,.$
Straightforward calculations show that $\kappa^{*}(0)=0<r$. Therefore, the Slater’s condition is satisfied, and by invoking strong duality we can write
$\sup_{w\in\mathbb{V}^{*}\colon\kappa^{*}(w)\leq r}{\langle w,f\rangle}$ $=\sup_{w\in\mathbb{V}^{*}}\inf_{\gamma\geq 0}{\langle w,f\rangle}-\gamma(\kappa^{*}(w)-r)$
$=\inf_{\gamma\geq 0}\sup_{w\in\mathbb{V}^{*}}{\langle w,f\rangle}-\gamma(\kappa^{*}(w)-r)$
$=\inf_{\gamma\geq 0}\gamma\kappa(\gamma^{-1}f)+\gamma r$
$=T_{r}(f)\,,$
where the last equation follows by the change of variable $\lambda=1/\gamma$. ∎
Proof of Lemma 2.
By the standard Chernoff bound, for any $T>T_{r}(f)$ we have
$\operatorname{\mathbb{P}}\left(\operatorname{\mathbb{E}}_{n}f(X)>T\right)$ $\leq\inf_{\lambda\geq 0}e^{n\log\operatorname{\mathbb{E}}e^{\lambda f(X)}-n\lambda T}\,.$
It follows from the definition of $T_{r}(\cdot)$ that there exists $\lambda^{\prime}\geq 0$ such that
$T_{r}(f)\leq\frac{r+\log\operatorname{\mathbb{E}}e^{\lambda^{\prime}f(X)}}{\lambda^{\prime}}<T\,.$
Therefore, we deduce
$\operatorname{\mathbb{P}}\left(\operatorname{\mathbb{E}}_{n}f(X)>T\right)$ $\leq e^{n\log\operatorname{\mathbb{E}}e^{\lambda^{\prime}f(X)}-n\lambda^{\prime}T}$
$\leq e^{-nr}\,.$
and consequently
$\operatorname{\mathbb{P}}\left(\operatorname{\mathbb{E}}_{n}f(X)\leq T_{r}(f)\right)$ $=\lim_{T\downarrow T_{r}(f)}\operatorname{\mathbb{P}}\left(\operatorname{\mathbb{E}}_{n}f(X)\leq T\right)$
$\geq 1-e^{-nr}\,.$